5/9 = 4/4x-3
9*4=36
5*2x-3=10x-3=7x
7x/36 <- Is that right?

Answer by Alan3354(69443)

5/9 = 4/4x-3
9*4=36
5*2x-3=10x-3=7x ************** 5*(2x-3) = 10x-15 but it's (4x-3)
==============================
7x/36 <- Is that right? Not right.